. You have drawn a simple random sample of 100 sociology majors for your senior research project. When reporting your sample data, can you also estimate population parameters? Explain/justify your answer.
Answer #1

We can estimate the population parameters as follows:

n = Sample Size

Let xi , (i = 1 to 500) be the values of the sample of 500 sociology majors of our senior research project.

Calculate:

Sample mean (\bar{x})

and

Sample Standard Deviation (s)

from the sample values.

The point estimate of population mean (\sigma ) is the sample mean (\bar{x} ).

The confidence Interval for population mean (\sigma ) is given by:

(\bar{x} )\pm Z_{\alpha /2} \times SE,

where

\alpha is significance level and SE is the Standard Error.